if you want to hit a baseball in such a way that the ball goes as far as possible, what angle of inclination should you aim for? Explain.
Answer #1

Range of projectile is given by:

Range = V^2* sin (2thetha) / g

It will be maximum if  sin (2thetha) = 1

sin (2thetha) = 1

2thetha = 90 degree

thetha = 90/2 = 45 degree

Answer:45 degree